OpenStudy (anonymous):

find the constant a , such that the function is continuous on the entire real number line. g(x) = x^2 - a^2/ x-a , where the x is not equal to a 8 , where x is equal to a. so, how can i know the contanst of a ? please help me

OpenStudy (anonymous):

\(a\) is not equal to \(8\) ?

OpenStudy (anonymous):

no matter, here is the gimmick factor and cancel , then replace \(x\) by \(a\)

OpenStudy (anonymous):

\[\frac{x^2-a^2}{x-a}=\frac{(x+a)(x-a)}{x-a}=x+a\] if \(x\neq a\)

OpenStudy (anonymous):

the limit as \(x\to a\) will be \(2a\) question is worded strangely

OpenStudy (anonymous):

ooooh i see !! it is find \(a\) so that the limit is \(8\)! i get it

OpenStudy (anonymous):

since as we see above, the limit will be \(2a\) so it \(2a=8\) then \(a=4\) doe, sorry i didn't understand the question

OpenStudy (anonymous):

@satellite73 thanks for the response =) but how u got the 2a? it's a bit confusing. hee

hartnn (hartnn):

\(\Large \lim \limits_{x \to a}f(x) = f(a)\) basically, whenever you can, plug in x=a directly in the function.

hartnn (hartnn):

\(\Large x+a |_{x \to a} = a+a = 2a\)

OpenStudy (anonymous):

ok thanks @hartnn . i got it =)

hartnn (hartnn):

welcome ^_^